How many liters of 70 percent alcohol must be mixed with 150 liters of 10 percent alcohol to make a mixture that is 50 percent alcohol
Answer by richwmiller(17219)

We have 150 liter of 10% solution
We want to add some 70% solution to make 50% solution
0.7*x+0.1*150=0.5(150+x)
0.7*x+15=75+0.5x
0.7x-0.5x=75-15
0.2x=60
x=300 liter at 70% (210) is added to 150 liter at 10%(15) to make 50% of 450 liter(225)
check
0.7*300+0.1*150=0.5(150+300)
210+15=0.5(450)
225=225
ok
codemixadd